# CRVQ: Channel-relaxed Vector Quantization for Extreme Compression of LLMs

### v1

@@ 6.4 Combination with Stronger VQ Settings
Although traditional VQ can enhance its performance by increasing $e$, we recommend using $e=8$ in combination with additional extended codebooks when deploying models. First, increasing $e$ (e.g., to $e=10$) results in a larger overall bit-width, leading to lower efficient bit utilization. It also hampers hardware alignment and computational acceleration. Furthermore, combining hardware-friendly basic codebooks with several extended codebooks and adjusting $\lambda\%$ critical channels can achieve superior quantization performance without these drawbacks. From this perspective, CRVQ can be flexibly adapted to accommodate the resource of different devices.
6.5 Inference Speed
@@
6.6 Comparison to QAT method
Our work focuses on extreme algorithms within the PTQ area. A related direction with similar goal is QAT, and the two methods have their respective advantages and limitations. QAT achieves relatively better performance through extensive training, whereas PTQ avoids the long training process but often falls short of QAT in performance. Our study shows that with continued advancements, PTQ is approaching the performance levels of QAT. We compare our method with the classical 1-bit QAT approach, OneBit [37], as shown in Table 4. Notably, our algorithm outperforms OneBit in terms of computational efficiency while achieving comparable performance.
Table 4: Comparison between OneBit and CRVQ on LLaMA2-13B. We use $8+8\times 3$ setting.
Method Wiki2 #GPUs #Hours /GPU #Hours
OneBit 8.76 16 448 7168
CRVQ 9.81 1 50 50
7 Conclusion
We propose a simple yet effective extreme quantization algorithm CRVQ and provide detailed discussions. The core idea is to carefully identify critical weight channels and apply multi-codebook fitting to a small subset of weights, achieved through channel reordering. Extensive experiments on models of varying sizes across OPT and LLaMA demonstrate that CRVQ leads significant improvements in quantization performance at minimal additional cost, outperforming representative strong baselines. We further analyze the factors influencing the effectiveness of CRVQ and provide suggestions for flexibly adapting the method to hardware platforms.
Appendix A Appendix
A.1 Details on Baselines and CRVQ
@@
  • PB-LLM [40]: It is an early approach to sub-2-bit PTQ, introduces mixed-precision quantization by applying higher bit-widths to a small subset of sensitive weights while binarizing the majority. We quantize the majority to 1-bit and maintain the sensitive weights to 8bit, leading to a 1.7-bit algorithm. We use hessian metric to find the salient weights. The other parameters are based on the code released by the authors.
@@
  • BiLLM [21]: It enhances quantization precision by partitioning the weight distribution and assigning different quantization parameters to each partition. We use hessian metric to partition the weights.
@@
  • AQLM [14]: AQLM is a vector quantization method that combines layer-level, block-level, and end-to-end fine-tuning. We limit the fine-tuning of intra-layer codebooks to a maximum of 100 epochs, typically not exceeding 20 epochs in most cases. For block-level and end-to-end fine-tuning, we perform a maximum of 25 epochs, using a batch size of 64. The fine-tuning process are optimized by Adam with $\beta_{1}=0.90,\beta_{2}=0.95$.
@@
  • CRVQ: The fine-tuning parameters in our method are the same as those used in AQLM. Given the use of multiple codebooks, we set the beam search width [14] to 1.

### v2

@@ 6.5 Hardware-friendly Deployment
Although traditional VQ can also enhance its performance by increasing $e$, we recommend using $e=8$ in combination with additional extended codebooks when deploying models. First, increasing $e$ (e.g., to $e=10$) may result in a larger overall bit-width, leading to lower bit-width utilization. It also hampers hardware alignment and computational acceleration. Furthermore, combining hardware-friendly basic codebooks with several extended codebooks and adjusting $\lambda\%$ critical channels can efficiently achieve better quantization performance without these drawbacks. Therefore, CRVQ is cost-effective on low-resource devices and easily adaptable to different devices.
6.6 Comparison to QAT method
Our work focuses on extreme algorithms within the PTQ area. A related direction with similar goal is QAT, and the two methods have their respective advantages and limitations. QAT achieves relatively better performance through extensive training, whereas PTQ avoids the long training process but often falls short of QAT in performance. We compare our method with the classical 1-bit QAT approach, OneBit (Xu et al., 2024), as shown in Table 3. Notably, our algorithm outperforms OneBit in terms of computational efficiency while achieving comparable performance. Our study shows that with continued advancements, PTQ is approaching the performance levels of QAT.
Method Wiki2 #GPUs #Hours /GPU #Hours
OneBit 8.76 16 448 7168
CRVQ 9.81 1 50 50
Table 3: Comparison between OneBit and CRVQ on LLaMA2-13B. Here we use $8+8\times 3$ setting.
7 Conclusion
We introduce CRVQ, an simple yet effective extreme quantization algorithm that enhances performance by identifying critical weight channels and applying multi-codebook fitting via channel reordering. Extensive experiments on models of varying sizes demonstrate significant performance gains at minimal additional cost, outperforming representative strong baselines. We also analyze key factors affecting CRVQ and provide guidance for hardware resource adaptation.
Limitations
While our method achieves obvious post-training extreme compression with minimal cost, offering a cost-effective solution for deploying LLMs on low-resource devices, some limitations remain. First, like other 1-bit baselines (Xu et al., 2024; Huang et al., 2024), our approach exhibits performance degradation compared to the original model and higher-bit quantization methods (e.g., 2-bit and above), highlighting the need for further exploration into the limits of extreme compression. Additionally, although our method significantly reduces computational demands compared to QAT-based 1-bit quantization (Xu et al., 2024), the quantization process itself remains time-consuming. Fortunately, this is a one-time cost. Finally, similar to many classic quantization studies, the role of a small subset of critical channels in the quantization process and its impact on overall performance remain insufficiently explored. This opens up exciting avenues for future research.
Ethics Statement
In this study, we employ models that are publicly available and open source. We affirm that the use of these models aligns with their original intended purposes. These models have been utilized strictly within the scope of academic and research-based activities, adhering to ethical guidelines and ensuring compliance with open-source licenses.
Appendix A Appendix
A.1 Average Bits Calculation
We first compute the average bit-width of quantized weights using the classical vector quantization method and explain the derivation of Equation 6, with a 4096$\times$4096 matrix as an example. Following the notations in this paper, let $m$ denote the number of codebooks, $e$ the codebook bit-width, $d$ the dimension of the short vectors obtained from matrix partitioning, and $n$ the average bit-width. The total space consumption consists of two parts: the codebooks and the binary encodings.
Consider the total number of vectors. Each row is divided into $2^{12}/d$ vectors, and the entire matrix is partitioned into $2^{24}/d$ vectors. If each vector is represented using $e$-bit binary encoding, the total code space required is $2^{24}e/d$ bits. Since each codebook has its own code, the total code overhead is $2^{24}me/d$ bits. Additionally, each codebook contains $2^{e}$ vectors of length $d$, where each value is stored in FP16 format. Therefore, the total space occupied by $m$ codebooks is $2^{e}md\times 2^{4}$ bits. Thus, the total storage requirement is given by $2^{24}me/d+2^{e+4}md$. Since the matrix contains $2^{24}$ elements, the quantized average bit-width is:
$n=\frac{me}{d}+2^{e-20}md.$
Assuming that the setting of AQLM is $m=1,e=8,d=8$, the average bit-width here is $n\approx 1.002$.
CRVQ introduces the concept of extended codebooks. Suppose that among the $m$ codebooks, there are $m-1$ extended codebooks used to encode $\lambda$-fraction of important vectors. The total code space for these important vectors can be expressed as $(m-1)\lambda e/d$. Moreover, storing column reordering indices requires extra space. Each index occupies 16 bits (with 12 effective bits), resulting in an additional storage requirement of $2^{12}\times 2^{4}/2^{24}$ bits. Clearly, this overhead is negligible. Thus, the total storage requirement for CRVQ is given by:
$n=\frac{e}{d}+\frac{(m-1)\lambda e}{d}+2^{e-20}md.$ (8)
Assuming that the setting of CRVQ is $8+8\times 3$ (i.e., $m=4$), $d=8$ and $\lambda=0.02$, the average bit-width here is $n\approx 1.062$.
A.2 Hessian-based Importance Metric
As shown in Equation 7, the importance metric based on the Hessian proxy considers both weight quantization error and activation magnitude. To understand the computation of importance, we unpack the matrix formulation. Let $\mathbf{W}^{M\times N}$ denote the weight matrix to be quantized, and the quantization error for each weight can be represented by $\Delta w_{ji}=w_{ji}-\mathrm{VQ}(w_{ji})$. Suppose the activation multiplying with the weight is denoted as $\mathbf{X}^{N\times O}$. Then, the Hessian proxy can be computed as:
$\mathbf{XX}^{\mathrm{T}}=\begin{pmatrix}\sum_{k=1}^{O}x_{1k}^{2}&\cdots&\sum_{k=1}^{O}x_{1k}x_{Nk}\\ \vdots&\ddots&\vdots\\ \sum_{k=1}^{O}x_{Nk}x_{1k}&\cdots&\sum_{k=1}^{O}x_{Nk}^{2}\end{pmatrix}$ (9)
Hence the Hessian-based importance metric can be reformulated as:
$I_{i}=\max_{j}\frac{1}{2}\sum_{k=1}^{O}(x_{ik}\Delta w_{ji})^{2}.$ (10)
Therefore, the selection of critical channels is guided by the outcome of the matrix product. Channels with both high quantization error and large corresponding activation magnitude have the greatest impact on the quantization results.
A.3 Results on LLaMA3 and Qwen2
To further validate the effectiveness of CRVQ across more LLMs, we extend our experiments to LLaMA3 and Qwen2 models, ranging in size from 0.5B to 8B. Except for the differences in model series, all experimental settings remain consistent with those in Section 5.1, including the use of Red_Pajama as the calibration dataset. Additionally, the evaluation datasets follow the same configuration as described in Section 5.1.
Here, we compare CRVQ against two strong baselines, BiLLM and AQLM, with results presented in Table A.3. The experiments show that, in terms of perplexity, CRVQ consistently outperforms all baselines across different models. For zero-shot accuracy, CRVQ surpasses other baselines on most datasets, achieving overall superior performance.
Model Method Wbits Perplexity($\downarrow$) Zero-shot Accuracy($\uparrow$)
Wiki2 C4 Wino. Hella. PIQA BoolQ ARC-e ARC-c Avg.
LLaMA3.2-1B FP16 16.0 9.07 12.04 60.14 63.62 74.59 63.98 60.52 36.01 59.81
[…]
AQLM 1.01 115.25 90.85 47.91 26.95 53.92 50.00 30.85 20.65 38.38
CRVQ 1.06 70.03 64.72 48.70 28.15 54.95 61.47 31.19 21.59 41.01
LLaMA3.2-3B FP16 16.0 7.28 10.01 69.38 73.61 77.42 72.94 71.59 45.90 68.47
[…]
AQLM 1.01 83.03 63.99 50.04 28.16 53.97 61.62 30.89 21.16 40.97
CRVQ 1.08 49.17 40.23 50.91 30.98 56.04 62.45 32.03 21.42 42.31
LLaMA3.1-8B FP16 16.0 5.84 8.43 73.48 78.85 81.18 82.02 81.19 53.50 75.03
[…]
AQLM 1.01 57.90 48.44 49.57 28.77 54.19 57.71 30.13 20.22 40.09
CRVQ 1.07 27.36 29.53 51.62 36.10 59.92 61.53 37.25 23.29 44.95
Qwen2-0.5B FP16 16.0 12.35 16.18 57.46 49.07 69.31 60.86 50.25 28.67 52.60
[…]
AQLM 1.01 96.34 84.86 48.54 27.26 52.18 54.68 31.06 21.33 39.17
CRVQ 1.06 77.72 69.06 50.59 26.87 54.46 61.28 31.36 22.78 41.22
Qwen2-1.5B FP16 16.0 8.87 12.20 66.38 65.37 75.35 72.42 60.77 36.26 62.75
[…]
AQLM 1.01 52.24 51.97 53.35 29.10 56.26 62.05 31.02 21.59 42.22
CRVQ 1.07 44.62 43.66 51.70 29.68 56.47 62.23 31.40 22.18 42.28
Qwen2-7B FP16 16.0 6.67 9.57 72.45 78.85 81.01 84.98 74.49 49.91 73.61
[…]
AQLM 1.01 20.57 23.03 52.88 42.30 62.46 62.02 41.62 25.34 47.77
CRVQ 1.06 17.39 19.91 57.54 47.18 65.13 64.04 46.25 27.56 51.28
Table 4: Perplexity and zero-shot accuracy on LLaMA3 on Qwen2 models. The results of AQLM and CRVQ are based on the models after e2e-finetuning. The best scores are highlighted in bold.
Model Method Wbits Perplexity($\downarrow$) Zero-shot Accuracy($\uparrow$)
Wiki2 C4 Wino. Hella. PIQA BoolQ ARC-e ARC-c Avg.
Mistral-7B FP16 16.0 4.91 7.42 74.11 81.11 82.15 83.76 79.59 54.18 75.81
[…]
CRVQ 1.06 13.01 16.23 54.70 44.11 60.83 62.75 42.26 25.51 48.36
Mistral-8$\times$7B FP16 16.0 3.59 6.52 76.64 83.94 83.35 85.41 83.42 59.81 78.76
[…]
CRVQ 1.08 9.28 12.32 60.69 63.60 73.56 72.32 57.53 34.47 60.36
Table 5: Perplexity and zero-shot accuracy on Mistral-7B / 8$\times$7B models. The results of AQLM and CRVQ are based on the models after e2e-finetuning. The best scores are highlighted in bold.
A.4 Results on Mistral-7B / 8$\times$7B
We also extend the application of CRVQ to the Mistral model [^2] https://mistral.ai/en/models. Furthermore, to evaluate the effectiveness of our algorithm on larger model and MoE architectures, we conduct experiments on Mistral-8$\times$7B. All experimental settings and datasets remain consistent with those in Section 5.1.
We compare CRVQ with AQLM. Experimental results demonstrate that, compared to the strong baseline AQLM, CRVQ consistently achieves superior perplexity and downstream task accuracy. This conclusion holds even for larger MoE models.
Figure 6: Bit scaling laws of AQLM and CRVQ on perplexity and zero-shot accuracy (PIQA). In (a) and (b), we increase the AQLM codebook bit-width from 8 to 12 and compare it with CRVQ in different $\lambda$. Note that the basic codebook bit-width of CRVQ constantly equals 8 here. The hardware-unfriendly settings are noted in red. In (c), the codebook bit-width of AQLM and basic codebook bit-width of CRVQ synchronously increase from 8 to 12.
(a) vector samples from product of important weight channels and activation
(b) vector samples from product of non-important weight channels and activation
Figure 7: 2-D visualization of the element-wise product from quantized weight and activation. The weight matrix is q_proj from layer 0 of LLaMA2-7B and we partition the product into vectors with a length of $d=8$. Here we use $\lambda=0.02$. (a) corresponds to important channels, while (b) corresponds to non-important channels. We compare the impact of different numbers of extended codebooks (ranging from 0 to 3) and channel importance on the distribution of product vectors.
A.5 Bit Scaling Laws
In Section 6.5, we discuss the device flexibility of CRVQ. To further illustrate its bit scaling property and deployment efficiency, we evaluate its perplexity and zero-shot accuracy under different bit-width configurations.
Compared to previous works, both AQLM and CRVQ allow flexible bit-width customization, enabling models to scale their capabilities according to hardware constraints. However, in extreme compression scenarios (sub-2-bit), AQLM can only adjust the codebook bit-width $e$ to control bit-width, as increasing the number of codebooks $m$ would double the total bit-width, exceeding 2. Yet, setting $e$ to 9, 10, 11, or 12 is impractical for deployment, as it misaligns memory boundary with hardware-friendly configurations, leading to inefficiencies and limited acceleration (Egiazarian et al., 2024). This approach, even if effective, is not practical for real deployment (see Figure 6(a) and 6(b) for details).
In contrast, CRVQ offers a more deployment-friendly solution by keeping the basic codebook bit-width fixed at $e=8$ and adjusting the critical channel ratio $\lambda$ to achieve bit scaling, thereby enabling capability scaling. This method achieves similar effectiveness while maintaining a hardware-efficient computation process.
Additionally, if not considering hardware efficiency, we investigate how increasing the basic codebook bit-width in CRVQ affects its advantage over AQLM. In practice, as the basic codebook bit-width increases and model capacity improves, the advantage over AQLM gradually diminishes. This trend is illustrated in Figure 6(c). The reason for this lies in the application scenario of CRVQ, which focuses on exhuming bit-width efficiency in extreme compression. When the basic codebook provides sufficient ability, the contribution of critical channels becomes negligible.
A.6 Codebook Fitting Visualization
We have demonstrated that a small subset of critical channels play a significant role in vector quantization. In our previous analysis, this is primarily reflected in overall LLM perplexity and downstream task accuracy. However, we remain curious about the differences between applying extended codebooks to important versus non-important channels. To figuratively understand this distinction, we partition the element-wise product of the quantized weight matrix and an activation vector into vectors of $d=8$, then apply PCA for 2-D visualization. The results from important and non-important channels are shown in Figure 7(a) and 7(b), respectively.
From the Figure 7, it is evident that the product vectors from important channels and non-important channels exhibit distinct distribution patterns after dimensionality reduction. The former resembles a scattered point cloud, while the latter forms a cross-like shape. As the number of extended codebooks increases, the product vectors of important channels gradually transition from a disordered spread to a cross-like structure, accompanied by a subtle rotational effect. In contrast, the product vectors of non-important channels primarily exhibit rotation without significant changes in shape as the number of extended codebooks increases.
These observations suggest that differences in channel importance manifest different quantization behaviors. Notably, the important weights change significantly with the increase in the number of extended codebooks.
Model ARB-LLM CRVQ
Wbits Wiki2 Wbits Wiki2
LLaMA-7B 1.09 14.03 1.07 13.68
LLaMA-13B 1.09 10.18 1.06 10.73
LLaMA2-7B 1.08 16.44 1.07 13.01
LLaMA2-13B 1.08 11.85 1.06 9.81
LLaMA3.1-8B 1.06 27.42 1.07 27.36
Table 6: Comparison between ARB-LLM and CRVQ.
A.7 Comparison with ARB-LLM
[…]
@@ A.7 Comparison with ARB-LLM
A.8 Inference Speed
A.9 Details on Baselines and CRVQ
@@
  • PB-LLM (Yuan et al., 2024): PB-LLM, which can be integrated with other classical PTQ algorithms, is an early approach to sub-2-bit PTQ, introduces mixed-precision quantization by applying higher bit-widths to a small subset of sensitive weights while binarizing the majority. We quantize the majority to 1-bit and maintain the sensitive weights to 8-bit, leading to a 1.7-bit algorithm. We use hessian metric to find the salient weights. The other parameters are based on the code released by the authors.
@@
  • BiLLM (Huang et al., 2024): BiLLM enhances quantization precision by partitioning the weight distribution and assigning different quantization parameters to each partition. It also employs residual-based multi-step quantization to improve accuracy. We use hessian metric to partition the weights and other settings are identical to the original paper.
@@
  • AQLM (Egiazarian et al., 2024): AQLM is an additive vector quantization method that combines layer-level, block-level, and end-to-end fine-tuning. We limit the fine-tuning of intra-layer codebooks to a maximum of 100 epochs, typically not exceeding 20 epochs in most cases. For block-level and end-to-end fine-tuning, we perform a maximum of 25 epochs, using a batch size of 64. The fine-tuning process are optimized by Adam with $\beta_{1}=0.90,\beta_{2}=0.95$.
@@
  • CRVQ: The fine-tuning parameters in our method are the same as those used in AQLM. Given the use of multiple codebooks, we set the beam search width (Egiazarian et al., 2024) to 1, which is the same as the project AQLM (Egiazarian et al., 2024).
A.10 Supplemental Related Work
Model quantization research includes several subfields. Based on the quantization object, it can be categorized into weight-only quantization (Huang et al., 2024; Tseng et al., 2024; Frantar et al., 2022) and weight-activation quantization (Xiao et al., 2023; Dettmers et al., 2022). Weight-only quantization (WxA16) focuses on low-bit weight representation. In contrast, weight-activation quantization (WxAy) quantizes both weights and intermediate activations. Model quantization can also be classified into quantization-aware training (QAT) and post-training quantization (PTQ) based on its integrating stage.
As introduced in Section 2, this paper focuses on PTQ, and it also belongs to weight-only quantization. Due to space limitations, the survey and comparison of QAT are provided in the appendix.
QAT incorporates quantization into the training process, minimizing performance loss due to low precision and achieving superior results. LLM-QAT (Liu et al., 2024) introduces a few learnable parameters and employs knowledge distillation to transfer abilities from the original model. QLoRA (Dettmers et al., 2023) mitigates the adverse effects of quantization by fine-tuning high-precision low-rank adapters. Recently, researchers explore QAT for more effective extreme compression. EfficientQAT (Chen et al., 2024) adjusts all model through a novel block-wise fine-tuning, achieving outstanding results at the 2-bit quantization level while reducing computational overhead. OneBit (Xu et al., 2024) designs a new architecture for 1-bit linear layers, incorporating two high-precision adapter vectors and leveraging knowledge distillation to enable effective 1-bit quantization. BinaryMoS (Jo et al., 2024) further improves on OneBit by replacing the vectors with a mixture-of-experts paradigm. Although QAT is effective for extreme quantization, it still faces challenges such as performance constraints and substantial computational costs. We provide comparison between our CRVQ (PTQ) and OneBit (QAT) in Section 6.6.
A.11 Generation Cases
We quantize the LLaMA2-13B model using different methods and compare their generated outputs for given prompts, as shown in Table 8. The average bit-width of the models is 1.08, 1.01, and 1.06 bits respectively. The results indicate that CRVQ produces longer and more fluent text. However, despite its fluency, the generated content still contains many hallucinations. For example, in case 3, the capital of New York State is 150 miles from New York City, not 1,100 miles, and there are 62 counties, not 10.
Prompt A cowboy rides a_
BiLLM cowboy rides a cowboy
AQLM troller. in the troller. The troller is a troller.
CRVQ 1950s Ford Mustang in the desert. The cowboy rides a 1950s Ford Mustang in the desert.
Prompt There are a number of ways to reduce air pollution, such as_
BiLLM 100% of the air pollution. The air pollution is a number of the air pollution.
AQLM 10000000000000000000000
CRVQ 1) reducing the amount of pollution from the burning of fossil fuels, 2) reducing the amount of pollution from the burning of coal, 3) reducing the amount of pollution from the burning of natural gas, 4) reducing the amount of pollution from the burning of wood, 5) reducing the amount of pollution from the burning of natural gas
Prompt The capital of the state of New York is_
BiLLM 100% of the state of New York is 100% of the state of New York
AQLM 100% of the state of New York. of the state of New York
CRVQ 1100 miles away from the city of New York. The state of New York is divided into 10 counties.
Table 8: Generation cases of CRVQ and baselines. The three prompt we use is from (Jo et al., 2024).